### v1

# A Note on Estimation Error Bound and Grouping Effect of Transfer Elastic Net

Abstract
The Transfer Elastic Net is an estimation method for linear regression models that combines $\ell_{1}$ and $\ell_{2}$ norm penalties to facilitate knowledge transfer. In this study, we derive a non-asymptotic $\ell_{2}$ norm estimation error bound for the estimator and discuss scenarios where the Transfer Elastic Net effectively works. Furthermore, we examine situations where it exhibits the grouping effect, which states that the estimates corresponding to highly correlated predictors have a small difference.
Keywords: High-dimensional data, Highly correlated variables, Linear regression model, Regularized estimation, Restricted eigenvalue condition, Sparse estimation, Transfer learning
1 Introduction
Regularized estimation of regression models involves estimation methods to achieve specific desirable properties in the resulting estimators through penalization terms added to the ordinary loss function. The Least Absolute Shrinkage and Selection Operator (Lasso) promotes sparsity in the coefficient estimates by applying a penalty based on the $\ell_{1}$ norm of the parameters, leading to some coefficients being exactly zero (Tibshirani,, 1996). Similarly, the Elastic Net encourages sparse estimates via the combined penalty of the $\ell_{1}$ norm and $\ell_{2}$ norm (Zou and Hastie,, 2005). Thanks to the $\ell_{2}$ norm penalty, the Elastic Net exhibits the grouping effect, wherein highly correlated predictors tend to be assigned coefficient estimates with small differences (Zou and Hastie,, 2005; Zhou,, 2013). Therefore, the Elastic Net effectively addresses correlations among predictors.
Transfer learning refers to a set of techniques to utilize the knowledge acquired from solving source problems to solve another target problem. Recently, Takada and Fujisawa, (2020) proposed the Transfer Lasso, which uses the $\ell_{1}$ norm to transfer the information of the source estimates to the target problem, which encourages the sparsity in the estimates and the sparsity in the changes of the estimates. Tomo and Nakaki, (2024) proposed the Transfer Elastic Net to address high-dimensional features in the field of bioinformatics, which employs a combination of $\ell_{1}$ and $\ell_{2}$ norms to facilitate knowledge transfer. Although Colangelo, (2021) proposed a similar estimator, they do not use their $\ell_{2}$ norm for knowledge transfer. In the following section, we define the estimator of the Transfer Elastic Net proposed in Tomo and Nakaki, (2024).
Suppose that we have responses $y_{i}\in\mathbb{R}$ ($i=1,\ldots,n$) and the predictors $\boldsymbol{X}_{i}=(X_{i,0},\ldots,X_{i,p-1})^{\top}\in\mathbb{R}^{p}$ ($i=1,\ldots,n$) corresponding to each response. The responses are centered as $\sum_{i=1}^{n}y_{i}=0$ and the predictors are standardized such that $\sum_{i=1}^{n}X_{i,j}=0$ and $\sum_{i=1}^{n}X_{i,j}^{2}/n=1$ for $j=0,\ldots,p-1$. We assume that the responses are generated according to a linear regression model with coefficients parameter ${\boldsymbol{\beta}}=({\beta}_{0},\ldots,{\beta}_{p-1})^{\top}\in\mathbb{R}^{p}$. Specifically, let $\boldsymbol{\beta}^{*}\in\mathbb{R}^{p}$ denote the true parameter vector, and we assume the responses are generated as
$y_{i}={\boldsymbol{\beta}^{*}}^{\top}\boldsymbol{X}_{i}+\varepsilon_{i},\quad\text{($i=1,\ldots,n$)},$
where $\varepsilon_{i}\in\mathbb{R}$ ($i=1,\ldots,n$) are the error terms with their expected values of 00. Assume that $\boldsymbol{\beta}^{*}\in\mathbb{R}^{p}$ has exactly $s$ non-zero elements ($0\leq s\leq p$). Let $\boldsymbol{y}=\left(y_{1},\ldots,y_{n}\right)^{\top}\in\mathbb{R}^{n}$ and $\mathbf{X}=(\boldsymbol{X}_{1},\ldots,\boldsymbol{X}_{n})^{\top}\in\mathbb{R}^{n\times p}$. Let $\tilde{\boldsymbol{\beta}}\in\mathbb{R}^{p}$ denote the parameter estimate obtained from the source problem. Then, the loss function of the Transfer Elastic Net is defined as:
$\mathcal{L}(\boldsymbol{\beta};\tilde{\boldsymbol{\beta}})$ $:=\frac{1}{2n}\sum_{i=1}^{n}\left(y_{i}-\boldsymbol{\beta}^{\top}\boldsymbol{X}_{i}\right)^{2}+\lambda R(\boldsymbol{\beta},\tilde{\boldsymbol{\beta}};\alpha,\rho),$ (1)
$R(\boldsymbol{\beta},\tilde{\boldsymbol{\beta}};\alpha,\rho)$ $:=\alpha\left\{\rho\|\boldsymbol{\beta}\|_{1}+(1-\rho)\|\boldsymbol{\beta}\|_{2}^{2}\right\}+(1-\alpha)\left\{\rho\|\boldsymbol{\beta}-\tilde{\boldsymbol{\beta}}\|_{1}+(1-\rho)\|\boldsymbol{\beta}-\tilde{\boldsymbol{\beta}}\|_{2}^{2}\right\},$
where $\lambda\in[0,+\infty)$, $\alpha\in[0,1]$, and $\rho\in[0,1]$ are the tuning parameters, which control the intensity of the regularization, the balance of the $\ell_{1}$ and $\ell_{2}$ norm, and the extent of transferring information, respectively. The estimator $\hat{\boldsymbol{\beta}}_{\mathrm{TENet}}$ is defined as the minimizer of the loss function:
$\hat{\boldsymbol{\beta}}_{\mathrm{TENet}}$ $:=\underset{\boldsymbol{\beta}\in\mathbb{R}^{p}}{\operatorname{argmin}}\mathcal{L}(\boldsymbol{\beta};\tilde{\boldsymbol{\beta}}).$
Takada and Fujisawa, (2020) introduced the generalized restricted eigenvalue condition and derived a non-asymptotic estimation error bound of the Transfer Lasso under the assumption of the sub-Gaussian error terms and the condition. In this study, we extend this analysis by establishing an estimation error bound for the Transfer Elastic Net. Then, we compare this bound with those of the ordinary Elastic Net and the Transfer Lasso to discuss the scenario where the Transfer Elastic Net is recommended to be applied. Furthermore, we derive an upper bound of the magnitude of the difference in estimates assigned to two different predictors and examine the situation where a high correlation between the predictors results in a small difference in the estimates.
2 Estimation Error Bound
Let $\boldsymbol{\Delta}=\tilde{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}$ and $\boldsymbol{\Delta}_{\alpha}=\boldsymbol{\Delta}-\alpha\tilde{\boldsymbol{\beta}}$ ($={(1-\alpha)\tilde{\boldsymbol{\beta}}-\boldsymbol{\beta}}^{*}$). Let $\boldsymbol{v}_{S}$ denote the subvector of $\boldsymbol{v}\in\mathbb{R}^{p}$ obtained by extracting the entries indexed by the index set $S$. We assume that $\boldsymbol{\beta}_{S}^{*}\in\mathbb{R}^{s}$ is non-zero in all its elements, and $\boldsymbol{\beta}_{S^{c}}^{*}\in\mathbb{R}^{p-s}$ is the zero vector. We establish the following assumptions to derive an estimation error bound for the Transfer Elastic Net.
Assumption 1 (Sub-Gaussianity of the error terms).
The error terms $\varepsilon_{1},\ldots,\varepsilon_{n}$ satisfies the conditions below.
  • i.
    $\varepsilon_{1},\ldots,\varepsilon_{n}$ are independent and identically distributed samples from a probability distribution $P_{\varepsilon}$ with its expected value of 00.
  • ii.
    $\varepsilon_{1},\ldots,\varepsilon_{n}$ are sub-Gaussian with variance proxy $\sigma^{2}$, i.e., for $i=1,\ldots,n$,
    $\mathrm{E}_{\varepsilon_{i}\sim P_{\varepsilon}}[\exp(t\varepsilon_{i})]$ $\leq\exp\left(\frac{\sigma^{2}t^{2}}{2}\right),\quad\forall t\in\mathbb{R}.$
Assumption 2 (Generalized restricted eigenvalue condition).
We define $\mathcal{B}(\alpha,\rho,c,\boldsymbol{\Delta})\subset\mathbb{R}^{p}$ as
$\mathcal{B}(\alpha,\rho,c,\boldsymbol{\Delta}):=\{\boldsymbol{v}\in\mathbb{R}^{p}$ $:(\alpha\rho-c)\|\boldsymbol{v}_{S^{c}}\|_{1}+(1-\alpha)\rho\|\boldsymbol{v}-\boldsymbol{\Delta}\|_{1}$
$\leq(\alpha\rho+c)\|\boldsymbol{v}_{S}\|_{1}+(1-\alpha)\rho\|\boldsymbol{\Delta}\|_{1}+2(1-\rho)\|\boldsymbol{\Delta}_{\alpha}\|_{2}\|\boldsymbol{v}\|_{2}\},$
for $\alpha\in[0,1]$, $\rho\in[0,1]$, $\boldsymbol{\Delta}\in\mathbb{R}^{p}$, and $c>0$. Then, the generalized restricted eigenvalue condition in Definition 1 of Takada and Fujisawa, (2020) holds for $\mathcal{B}=\mathcal{B}(\alpha,\rho,c,\boldsymbol{\Delta})$ for all $\alpha\in[0,1]$, $\rho\in[0,1]$, $\boldsymbol{\Delta}\in\mathbb{R}^{p}$, and $c>0$, i.e.,
$\phi(\mathcal{B}):=\inf_{\boldsymbol{v}\in\mathcal{B}}\frac{\boldsymbol{v}^{\top}\frac{1}{n}\mathbf{X}^{\top}\mathbf{X}\boldsymbol{v}}{\|\boldsymbol{v}\|_{2}^{2}}>0.$
Under these assumptions, we establish Theorem 1, which provides a non-asymptotic estimation error bound for the Transfer Elastic Net.
Theorem 1.
Suppose that Assumption 1 and Assumption 2 are satisfied. We define
$U_{\mathrm{TENet}}$ $:=\frac{(\alpha\rho+c)\lambda\sqrt{s}+2\lambda(1-\rho)\left\|\boldsymbol{\Delta}_{\alpha}\right\|_{2}+\sqrt{D}}{2\lambda(1-\rho)+\phi_{\mathrm{TENet}}},$
$D$ =$=\left\{(\alpha\rho+c)\lambda\sqrt{s}+2\lambda\alpha(1-\rho)\left\|\boldsymbol{\Delta}_{\alpha}\right\|_{2}\right\}^{2}$
$\quad+2\lambda(1-\alpha)\rho\|\boldsymbol{\Delta}\|_{1}\{2\lambda(1-\rho)+\phi_{\mathrm{TENet}}\},$
where $\phi_{\mathrm{TENet}}=\phi(\mathcal{B}(\alpha,\rho,c,\boldsymbol{\Delta}))$. Then we have
$\left\|\hat{\boldsymbol{\beta}}_{\mathrm{TENet}}-\boldsymbol{\beta}^{*}\right\|_{2}$ $\leq U_{\mathrm{TENet}},$
with probability at least $1-\exp(-nc^{2}\lambda^{2}/2\sigma^{2}+\log(2p))$.
Subsequently, we consider the ordinary Elastic Net and the Transfer Lasso. The estimator of the Elastic Net $\hat{\boldsymbol{\beta}}_{\mathrm{ENet}}$ is defined as
$\hat{\boldsymbol{\beta}}_{\mathrm{ENet}}:=\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{p}}\left\{\frac{1}{2n}\sum_{i=1}^{n}\left(y_{i}-\boldsymbol{\beta}^{\top}\boldsymbol{X}_{i}\right)^{2}+\lambda R(\boldsymbol{\beta},{\boldsymbol{0}};1,\rho)\right\},$
and, similarly, the estimator of the Transfer Lasso $\hat{\boldsymbol{\beta}}_{\mathrm{TLasso}}$ is defined as
$\hat{\boldsymbol{\beta}}_{\mathrm{TLasso}}:=\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{p}}\left\{\frac{1}{2n}\sum_{i=1}^{n}\left(y_{i}-\boldsymbol{\beta}^{\top}\boldsymbol{X}_{i}\right)^{2}+\lambda R(\boldsymbol{\beta},\tilde{\boldsymbol{\beta}};\alpha,1)\right\}.$
The estimation error for $\hat{\boldsymbol{\beta}}_{\mathrm{ENet}}$ and $\hat{\boldsymbol{\beta}}_{\mathrm{TLasso}}$ can be immediately obtained from Theorem 1 by setting $\alpha=1$ and $\rho=1$, respectively.
Corollary 2.
Suppose that Assumption 1 and Assumption 2 are satisfied. We define
$U_{\mathrm{ENet}}$ $:=\frac{2(\rho+c)\lambda\sqrt{s}+4\lambda(1-\rho)\left\|\boldsymbol{\beta}^{*}\right\|_{2}}{2\lambda(1-\rho)+\phi_{\mathrm{ENet}}},$
$U_{\mathrm{TLasso}}$ $:=\frac{(\alpha+c)\lambda\sqrt{s}+\sqrt{\left\{(\alpha+c)\lambda\sqrt{s}\right\}^{2}+2\lambda(1-\alpha)\|\boldsymbol{\Delta}\|_{1}\phi_{\mathrm{TLasso}}}}{\phi_{\mathrm{TLasso}}},$
where $\phi_{\mathrm{ENet}}=\phi(\mathcal{B}(1,\rho,c,\boldsymbol{0}))$ and $\phi_{\mathrm{TLasso}}=\phi(\mathcal{B}(\alpha,1,c,\boldsymbol{\Delta}))$. Then we have
$\left\|\hat{\boldsymbol{\beta}}_{\mathrm{ENet}}-\boldsymbol{\beta}^{*}\right\|_{2}$ $\leq U_{\mathrm{ENet}},$
and
$\left\|\hat{\boldsymbol{\beta}}_{\mathrm{TLasso}}-\boldsymbol{\beta}^{*}\right\|_{2}$ $\leq U_{\mathrm{TLasso}},$
with probability at least $1-\exp(-nc^{2}\lambda^{2}/2\sigma^{2}+\log(2p))$.
Remark 1.
The estimation error of Transfer Lasso described in Corollary 2 is consistent with the result of Takada and Fujisawa, (2020).
From Theorem 1 and Corollary 2, we establish the propositions that provide the comparison of the bounds. We assume that the source problem is highly related to the target problem and that $\tilde{\boldsymbol{\beta}}=\boldsymbol{\beta}^{*}$. Then, we prove the following proposition providing a relationship between the error bounds of the Transfer Elastic Net and the ordinary Elastic Net.
Proposition 3.
Suppose $\tilde{\boldsymbol{\beta}}=\boldsymbol{\beta}^{*}$. Then we have $U_{\mathrm{ENet}}\geq U_{\mathrm{TENet}}$.
We then consider the Transfer Lasso. The following proposition shows a sufficient condition under which $U_{\mathrm{TENet}}$ is guaranteed to be smaller than $U_{\mathrm{TLasso}}$.
Proposition 4.
Suppose $\tilde{\boldsymbol{\beta}}=\boldsymbol{\beta}^{*}$. If $\sqrt{s}/2\geq\|\boldsymbol{\beta}^{*}_{S}\|_{2}$ and $\phi_{\mathrm{TENet}}+2\lambda(1-\rho)\geq\phi_{\mathrm{TLasso}}$ , then we have $U_{\mathrm{TLasso}}\geq U_{\mathrm{TENet}}$.
These propositions suggest that the Transfer Elastic Net can achieve a lower estimation error bound than those of the Elastic Net and the Transfer Lasso when the source problem is highly related to the target problem. In particular, when the correlation between predictors is large, $\phi$ may take a value close to zero. Even in such a case, the error bound for the Transfer Elastic Net can be smaller than that of the Transfer Lasso if a sufficiently large $\lambda$ value is chosen.
3 Grouping Effect
The grouping effect of the Elastic Net ensures that the estimates of the coefficients corresponding to highly correlated predictors have a small difference (Zou and Hastie,, 2005; Zhou,, 2013). We show that the Transfer Elastic Net retains a similar property. Let $r_{jk}$ denote the correlation coefficient between the $j$th and $k$th column of $\mathbf{X}$. Then we prove the following theorem.
Theorem 5.
If $\hat{{\beta}}_{j}\hat{{\beta}}_{k}>0$ and $(\hat{{\beta}}_{j}-\tilde{{\beta}}_{j})(\hat{{\beta}}_{k}-\tilde{{\beta}}_{k})>0$, then, for $\rho\neq 1$ and $\lambda\neq 0$, we have
$\left|\hat{{\beta}}_{j}-\hat{{\beta}}_{k}\right|$ $\leq Z\sqrt{1-r_{jk}}+(1-\alpha)\left|\tilde{{\beta}}_{j}-\tilde{{\beta}}_{k}\right|,\quad Z=\sqrt{\frac{\left\|\boldsymbol{y}\right\|_{2}^{2}+2n\lambda(1-\alpha)\rho\|\tilde{\boldsymbol{\beta}}\|_{1}+2n\lambda(1-\alpha)(1-\rho)\|\tilde{\boldsymbol{\beta}}\|_{2}^{2}}{2n\lambda^{2}(1-\rho)^{2}}}.$
This theorem suggests that the estimates corresponding to the strongly correlated variables exhibit a small difference if the corresponding source estimates have a small difference.
Now suppose that the source estimate $\tilde{\boldsymbol{\beta}}$ was obtained by applying the ordinary Elastic Net to the data with the response ${\boldsymbol{y}^{\prime}}\in\mathbb{R}^{m}$ and the predictors ${\mathbf{X}}^{\prime}\in\mathbb{R}^{m\times p}$, with $\sum_{i=1}^{n}y^{\prime}_{i}=0$, $\sum_{i=1}^{n}X_{i,j}^{\prime}=0$, and $\sum_{i=1}^{n}{X_{i,j}^{\prime}}^{2}/n=1$ for $j=0,\ldots,p-1$. Therefore, we assume that $\tilde{\boldsymbol{\beta}}$ is defined as:
$\tilde{\boldsymbol{\beta}}=\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{p}}\left\{\frac{1}{2m}\sum_{i=1}^{m}\left(y_{i}^{\prime}-\boldsymbol{\beta}^{\top}\boldsymbol{X}^{\prime}_{i}\right)^{2}+\lambda^{\prime}R(\boldsymbol{\beta},{\boldsymbol{0}};1,\rho^{\prime})\right\}.$
Let $\tilde{r}_{jk}$ denote the correlation coefficient between the $j$th and $k$th column of $\mathbf{X}^{\prime}$. Then we have
$\left|\tilde{{\beta}}_{j}-\tilde{{\beta}}_{k}\right|$ $\leq\frac{\|\tilde{\boldsymbol{y}}\|_{2}\sqrt{2n(1-\tilde{r}_{jk})}}{2n\lambda^{\prime}(1-\rho^{\prime})},$
for $\rho^{\prime}\neq 1$ and $\lambda^{\prime}\neq 0$. This result suggests that, if there are highly correlated predictors in the source data and the corresponding predictors are also highly correlated in the target data, adopting Elastic Net estimates as the source estimates will ensure a small difference between the corresponding estimates of the Transfer Elastic Net.
Acknowledgement
The authors thank Takayuki Kawashima for valuable comments.
Appendix A Proof of Theorem 1
In this section, we write $\hat{\boldsymbol{\beta}}=\hat{\boldsymbol{\beta}}_{\mathrm{TENet}}$ and $\phi=\phi_{\mathrm{TENet}}$. First, we establish the following lemma.
Lemma 6.
Suppose that Assumption 1 is satisfied. Then, we have
$\frac{1}{n}\left({\mathbf{X}}^{\top}{\boldsymbol{\varepsilon}}\right)^{\top}\left(\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right)\leq c\lambda\left\|\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right\|_{1}\quad\text{for some $c>0$},$
with probability at least $1-\exp(-nc^{2}\lambda^{2}/2\sigma^{2}+\log(2p))$.
Proof.
From Hölder’s inequality, it follows that
$\frac{1}{n}\left({\mathbf{X}}^{\top}{\boldsymbol{\varepsilon}}\right)^{\top}\left(\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right)\leq\frac{1}{n}\left\|{\mathbf{X}}^{\top}{\boldsymbol{\varepsilon}}\right\|_{\infty}\left\|\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right\|_{1}.$
From Assumption 1, we have
$\mathrm{E}_{\varepsilon_{1},\ldots,\varepsilon_{n}\sim P_{\varepsilon}}\left[\exp\left(t\left(\frac{1}{n}{\mathbf{X}}^{\top}{\boldsymbol{\varepsilon}}\right)_{j}\right)\right]$ $=\mathrm{E}_{\varepsilon_{1},\ldots,\varepsilon_{n}\sim P_{\varepsilon}}\left[\exp\left(\frac{t}{n}{\sum_{i=1}^{n}{X_{i,j}}}{{\varepsilon}_{j}}\right)\right]$
$=\prod_{i=1}^{n}\mathrm{E}_{\varepsilon_{1},\ldots,\varepsilon_{n}\sim P_{\varepsilon}}\left[\exp\left(\frac{t}{n}{{X_{i,j}}}{{\varepsilon}_{j}}\right)\right]$
$\leq\prod_{i=1}^{n}\exp\left(\frac{\sigma^{2}\left(t{X_{i,j}/n}\right)^{2}}{2}\right)$
$=\exp\left(\frac{\left(\sigma^{2}\sum_{i=1}^{n}{X_{i,j}^{2}/n^{2}}\right)t^{2}}{2}\right)$
$=\exp\left(\frac{\left(\sigma/\sqrt{n}\right)^{2}t^{2}}{2}\right),\quad\forall t\in\mathbb{R}.$
Therefore, $\left(\frac{1}{n}{\mathbf{X}}^{\top}{\boldsymbol{\varepsilon}}\right)_{j}$ is sub-Gaussian with variance proxy $\left(\sigma/\sqrt{n}\right)^{2}$ and we then have
$P\left(\left|\left(\frac{1}{n}{\mathbf{X}}^{\top}{\boldsymbol{\varepsilon}}\right)_{j}\right|>t\right)$ $\leq 2\exp\left(-\frac{t^{2}}{2\left(\sigma/\sqrt{n}\right)^{2}}\right),\quad\forall t>0.$
From the above results, we have
$P\left(\left\|\frac{1}{n}{\mathbf{X}}^{\top}{\boldsymbol{\varepsilon}}\right\|_{\infty}>t\right)$ $=P\left(\bigcup_{j=0}^{p-1}\left|\left(\frac{1}{n}{\mathbf{X}}^{\top}{\boldsymbol{\varepsilon}}\right)_{j}\right|>t\right)$
$\leq\sum_{j=0}^{p-1}P\left(\left|\left(\frac{1}{n}{\mathbf{X}}^{\top}{\boldsymbol{\varepsilon}}\right)_{j}\right|>t\right)$
$\leq\sum_{j=0}^{p-1}2\exp\left(-\frac{t^{2}}{2\left(\sigma/\sqrt{n}\right)^{2}}\right)$
$=\exp\left(-\frac{nt^{2}}{2\sigma^{2}}+\log(2p)\right),\quad\forall t>0.$
Substituting $t$ for $c\lambda$ with some $c>0$ and subtracting both sides from $1$ completes the proof. ∎
As a preparation, the loss function of the Transfer Elastic Net (1) can be expressed as
$L(\boldsymbol{\beta},\tilde{\boldsymbol{\beta}})$ $=\frac{1}{2n}\|\boldsymbol{y}-\mathbf{X}\boldsymbol{\beta}\|_{2}^{2}+\lambda\alpha\rho\|\boldsymbol{\beta}\|_{1}+\lambda\alpha(1-\rho)\|\boldsymbol{\beta}\|_{2}^{2}+\lambda(1-\alpha)\rho\|\boldsymbol{\beta}-\tilde{\boldsymbol{\beta}}\|_{1}+\lambda(1-\alpha)(1-\rho)\|\boldsymbol{\beta}-\tilde{\boldsymbol{\beta}}\|_{2}^{2}$
$=\frac{1}{2n}\|\tilde{\boldsymbol{y}}-\widetilde{\mathbf{X}}\boldsymbol{\beta}\|_{2}^{2}+\lambda\alpha\rho\|\boldsymbol{\beta}\|_{1}+\lambda(1-\alpha)\rho\|\boldsymbol{\beta}-\tilde{\boldsymbol{\beta}}\|_{1}-2\lambda(1-\alpha)(1-\rho){\boldsymbol{\beta}}^{\top}\tilde{\boldsymbol{\beta}}+\lambda(1-\alpha)(1-\rho)\|\tilde{\boldsymbol{\beta}}\|_{2}^{2},$
where $\tilde{\boldsymbol{y}}=(\boldsymbol{y}^{\top},0)^{\top}\in\mathbb{R}^{n+p}$ and $\widetilde{\mathbf{X}}=(\mathbf{X}^{\top},\mathbf{I}\sqrt{2n\lambda(1-\rho)})^{\top}\in\mathbb{R}^{(n+p)\times p}$ with a p-dimensional zero-vector $\boldsymbol{0}$ and $p\times p$ identity matrix $\mathbf{I}$. Using this expression, we have
$L(\boldsymbol{\beta}^{*},\tilde{\boldsymbol{\beta}})$ $=\frac{1}{2n}\|\tilde{\boldsymbol{y}}-\widetilde{\mathbf{X}}\boldsymbol{\beta}^{*}\|_{2}^{2}+\lambda\alpha\rho\|\boldsymbol{\beta}^{*}\|_{1}+\lambda(1-\alpha)\rho\|\boldsymbol{\beta}^{*}-\tilde{\boldsymbol{\beta}}\|_{1}+\lambda(1-\alpha)(1-\rho)\|\boldsymbol{\beta}^{*}-\tilde{\boldsymbol{\beta}}\|_{2}^{2}$
$=\frac{1}{2n}\|\tilde{\boldsymbol{\varepsilon}}\|_{2}^{2}+\lambda\alpha\rho\|\boldsymbol{\beta}^{*}\|_{1}+\lambda(1-\alpha)\rho\|\boldsymbol{\beta}^{*}-\tilde{\boldsymbol{\beta}}\|_{1}-2\lambda(1-\alpha)(1-\rho){\boldsymbol{\beta}^{*}}^{\top}\tilde{\boldsymbol{\beta}}+\lambda(1-\alpha)(1-\rho)\|\tilde{\boldsymbol{\beta}}\|_{2}^{2},$
where $\tilde{\boldsymbol{\varepsilon}}=(\boldsymbol{\varepsilon}^{\top},-{\boldsymbol{\beta}^{*}}^{\top}\mathbf{I}\sqrt{2n\lambda(1-\rho)})^{\top}\in\mathbb{R}^{n+p}$ with $\boldsymbol{\varepsilon}=(\varepsilon_{1},\ldots,{\varepsilon}_{n})\in\mathbb{R}^{n}$.
Then, we prove Theorem 1 as follows.
Proof of Theorem 1.
From Lemma 6, for some $c>0$, we have
$\frac{1}{2n}\|\tilde{\boldsymbol{\varepsilon}}\|_{2}^{2}-\frac{1}{2n}\|\tilde{\boldsymbol{y}}-\widetilde{\mathbf{X}}\boldsymbol{\beta}^{*}\|_{2}^{2}$
$=\frac{1}{2n}\left\{\|\tilde{\boldsymbol{\varepsilon}}\|_{2}^{2}-\left\|\tilde{\boldsymbol{\varepsilon}}-\widetilde{\mathbf{X}}\left(\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right)\right\|_{2}^{2}\right\}$
$=\frac{1}{n}\left(\widetilde{\mathbf{X}}^{\top}\tilde{\boldsymbol{\varepsilon}}\right)^{\top}\left(\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right)-\frac{1}{2n}\left\|\widetilde{\mathbf{X}}\left(\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right)\right\|_{2}^{2}$
$=\frac{1}{n}\left({\mathbf{X}}^{\top}{\boldsymbol{\varepsilon}}\right)^{\top}\left(\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right)-2\lambda(1-\rho){\boldsymbol{\beta}^{*}}^{\top}\left(\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right)-\frac{1}{2n}\left\|\widetilde{\mathbf{X}}\left(\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right)\right\|_{2}^{2}$
$\leq c\lambda\left\|\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right\|_{1}-2\lambda(1-\rho){\boldsymbol{\beta}^{*}}^{\top}\left(\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right)-\frac{1}{2n}\left\|\widetilde{\mathbf{X}}\left(\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right)\right\|_{2}^{2},$
with probability at least $1-\exp(-nc^{2}\lambda^{2}/2\sigma^{2}+\log(2p))$. Therefore, it follows that
$L(\boldsymbol{\beta}^{*},\tilde{\boldsymbol{\beta}})-L(\hat{\boldsymbol{\beta}},\tilde{\boldsymbol{\beta}})$
$\leq c\lambda\left\|\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right\|_{1}-2\lambda(1-\rho){\boldsymbol{\beta}^{*}}^{\top}\left(\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right)-\frac{1}{2n}\left\|\widetilde{\mathbf{X}}\left(\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right)\right\|_{2}^{2}$
$\quad+\lambda\alpha\rho\|\boldsymbol{\beta}^{*}\|_{1}+\lambda(1-\alpha)\rho\|\boldsymbol{\beta}^{*}-\tilde{\boldsymbol{\beta}}\|_{1}-2\lambda(1-\alpha)(1-\rho){\boldsymbol{\beta}^{*}}^{\top}\tilde{\boldsymbol{\beta}}$
$\quad-\left\{\lambda\alpha\rho\|\hat{\boldsymbol{\beta}}\|_{1}+\lambda(1-\alpha)\rho\|\hat{\boldsymbol{\beta}}-\tilde{\boldsymbol{\beta}}\|_{1}-2\lambda(1-\alpha)(1-\rho){\hat{\boldsymbol{\beta}}}^{\top}\tilde{\boldsymbol{\beta}}\right\}$
$\leq c\lambda\left\|\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right\|_{1}-2\lambda(1-\rho){\left(\boldsymbol{\beta}^{*}-(1-\alpha)\tilde{\boldsymbol{\beta}}\right)}^{\top}\left(\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right)-\frac{1}{2n}\left\|\widetilde{\mathbf{X}}\left(\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right)\right\|_{2}^{2}$
$\quad+\lambda\alpha\rho\|\hat{\boldsymbol{\beta}}_{S}-\boldsymbol{\beta}^{*}_{S}\|_{1}-\lambda\alpha\rho\|\hat{\boldsymbol{\beta}}_{S^{c}}\|_{1}+\lambda(1-\alpha)\rho\|\boldsymbol{\beta}^{*}-\tilde{\boldsymbol{\beta}}\|_{1}-\lambda(1-\alpha)\rho\|\hat{\boldsymbol{\beta}}-\tilde{\boldsymbol{\beta}}\|_{1}$
[…]
$\leq c\lambda\left\|\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right\|_{1}+2\lambda(1-\rho){\left\|\boldsymbol{\beta}^{*}-(1-\alpha)\tilde{\boldsymbol{\beta}}\right\|_{2}}\left\|\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right\|_{2}-\frac{1}{2n}\left\|\widetilde{\mathbf{X}}\left(\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right)\right\|_{2}^{2}$
$\quad+\lambda\alpha\rho\|\hat{\boldsymbol{\beta}}_{S}-\boldsymbol{\beta}^{*}_{S}\|_{1}-\lambda\alpha\rho\|\hat{\boldsymbol{\beta}}_{S^{c}}\|_{1}+\lambda(1-\alpha)\rho\|\boldsymbol{\beta}^{*}-\tilde{\boldsymbol{\beta}}\|_{1}-\lambda(1-\alpha)\rho\|\hat{\boldsymbol{\beta}}-\tilde{\boldsymbol{\beta}}\|_{1}.$ (2)
From the definition of $\hat{\boldsymbol{\beta}}$, we have $L(\boldsymbol{\beta}^{*},\tilde{\boldsymbol{\beta}})-L(\hat{\boldsymbol{\beta}},\tilde{\boldsymbol{\beta}})\geq 0$. Therefore, when $\boldsymbol{v}=\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}$, we obtain
$c\lambda\left\|{\boldsymbol{v}}_{S}\right\|_{1}+c\lambda\left\|{\boldsymbol{v}}_{S^{c}}\right\|_{1}+2\lambda(1-\rho){\left\|\boldsymbol{\Delta}_{\alpha}\right\|_{2}}\left\|{\boldsymbol{v}}\right\|_{2}$
$\quad+\lambda\alpha\rho\|{\boldsymbol{v}}_{S}\|_{1}-\lambda\alpha\rho\|{\boldsymbol{v}}_{S^{c}}\|_{1}+\lambda(1-\alpha)\rho\|\boldsymbol{\Delta}\|_{1}-\lambda(1-\alpha)\rho\|{\boldsymbol{v}}-{\boldsymbol{\Delta}}\|_{1}$
$\quad\geq 0,$
i.e., $(\alpha\rho+c)\|\boldsymbol{v}_{S}\|_{1}+(1-\alpha)\rho\|\boldsymbol{\Delta}\|_{1}+2(1-\rho)\|\boldsymbol{\Delta}_{\alpha}\|_{2}\|\boldsymbol{v}\|_{2}$
$\geq(\alpha\rho-c)\|\boldsymbol{v}_{S^{c}}\|_{1}+(1-\alpha)\rho\|\boldsymbol{v}-\boldsymbol{\Delta}\|_{1}.$
This means that $\boldsymbol{v}=\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\in\mathcal{B}(\alpha,\rho,c,\boldsymbol{\Delta})$. From Assumption 2, it follows that
$\frac{1}{2n}\left\|\widetilde{\mathbf{X}}\left(\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right)\right\|_{2}^{2}$ $=\frac{1}{2n}\left(\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right)^{\top}\widetilde{\mathbf{X}}^{\top}\widetilde{\mathbf{X}}\left(\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right)$
$=\frac{1}{2}\frac{\left(\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right)^{\top}\left(\frac{\mathbf{X}^{\top}\mathbf{X}}{n}+2\lambda(1-\rho)\mathbf{I}\right)\left(\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right)}{\|\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\|_{2}^{2}}\|\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\|_{2}^{2}$
$\geq\frac{1}{2}\{2\lambda(1-\rho)+\phi\}\|\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\|_{2}^{2}.$
Therefore, from the inequality (Appendix A), we have
00 0 $\leq\lambda\left(\alpha\rho+c\right)\left\|\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right\|_{1}+2\lambda(1-\rho){\left\|\boldsymbol{\Delta}_{\alpha}\right\|_{2}}\left\|\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right\|_{2}-\frac{1}{2}\{2\lambda(1-\rho)+\phi\}\|\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\|_{2}^{2}$
$\quad+\lambda(1-\alpha)\rho\|\boldsymbol{\Delta}\|_{1}$
$\leq\left\{\lambda\left(\alpha\rho+c\right)\sqrt{s}+2\lambda(1-\rho){\left\|\boldsymbol{\Delta}_{\alpha}\right\|_{2}}\right\}\left\|\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right\|_{2}-\frac{1}{2}\{2\lambda(1-\rho)+\phi\}\left\|\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\right\|_{2}^{2}$
$\quad+\lambda(1-\alpha)\rho\|\boldsymbol{\Delta}\|_{1}.$
[…]
Solving this inequality with respect to $\|\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}\|_{2}$ completes the proof. ∎
Appendix B Proof of Proposition 3
As preparation, we prove the following lemma.
Lemma 7.
Suppose that Assumption 2 is satisfied. Then we have $\phi(\mathcal{B}(\alpha,\rho,c,\boldsymbol{0}))\geq\phi(\mathcal{B}(1,\rho,c,\boldsymbol{0}))$.
Proof.
$\mathcal{B}(\alpha,\rho,c,\boldsymbol{0})$ and $\mathcal{B}(1,\rho,c,\boldsymbol{0})$ are, respectively, expressed as
$\mathcal{B}(\alpha,\rho,c,\boldsymbol{0})$
=$=\{\boldsymbol{v}\in\mathbb{R}^{p}:(\alpha\rho-c)\|\boldsymbol{v}_{S^{c}}\|_{1}+(1-\alpha)\rho\|\boldsymbol{v}\|_{1}\leq(\alpha\rho+c)\|\boldsymbol{v}_{S}\|_{1}+2\alpha(1-\rho)\|\boldsymbol{\beta}^{*}\|_{2}\|\boldsymbol{v}\|_{2}\},$
and
$\mathcal{B}(1,\rho,c,\boldsymbol{0})$ =$=\{\boldsymbol{v}\in\mathbb{R}^{p}:(\rho-c)\|\boldsymbol{v}_{S^{c}}\|_{1}\leq(\rho+c)\|\boldsymbol{v}_{S}\|_{1}+2(1-\rho)\|\boldsymbol{\beta}^{*}\|_{2}\|\boldsymbol{v}\|_{2}\}.$
For the left side of the inequalities represented in the definition of $\mathcal{B}(\alpha,\rho,c,\boldsymbol{0})$ and $\mathcal{B}(1,\rho,c,\boldsymbol{0})$,
$(\rho-c)\|\boldsymbol{v}_{S^{c}}\|_{1}$ $=(\alpha\rho-c)\|\boldsymbol{v}_{S^{c}}\|_{1}+(1-\alpha)\rho\|\boldsymbol{v}_{S^{c}}\|_{1}$
$=(\alpha\rho-c)\|\boldsymbol{v}_{S^{c}}\|_{1}+(1-\alpha)\rho\|\boldsymbol{v}_{S^{c}}\|_{1}$
$\quad+(1-\alpha)\rho\|\boldsymbol{v}\|_{1}-(1-\alpha)\rho\|\boldsymbol{v}\|_{1}$
$=(\alpha\rho-c)\|\boldsymbol{v}_{S^{c}}\|_{1}+(1-\alpha)\rho\|\boldsymbol{v}\|_{1}$
$\quad-(1-\alpha)\rho\left(\|\boldsymbol{v}\|_{1}-\|\boldsymbol{v}_{S^{c}}\|_{1}\right)$
$\leq(\alpha\rho-c)\|\boldsymbol{v}_{S^{c}}\|_{1}+(1-\alpha)\rho\|\boldsymbol{v}\|_{1}.$
Subsequently, for the right side,
$(\alpha\rho+c)\|\boldsymbol{v}_{S}\|_{1}+2\alpha(1-\rho)\|\boldsymbol{\beta}^{*}\|_{2}\|\boldsymbol{v}\|_{2}$ $\leq(\rho+c)\|\boldsymbol{v}_{S}\|_{1}+2(1-\rho)\|\boldsymbol{\beta}^{*}\|_{2}\|\boldsymbol{v}\|_{2}.$
Therefore, we have $\mathcal{B}(\alpha,\rho,c,\boldsymbol{0})\subseteq\mathcal{B}(1,\rho,c,\boldsymbol{0})$ and then, $\phi(\mathcal{B}(\alpha,\rho,c,\boldsymbol{0}))\geq\phi(\mathcal{B}(1,\rho,c,\boldsymbol{0}))$. ∎
Based on this lemma, we establish the proof of Proposition 3 as follows.
Proof of Proposition 3.
Suppose $\tilde{\boldsymbol{\beta}}=\boldsymbol{\beta}^{*}$. Then, we have
$U_{\mathrm{TENet}}$ $=\frac{2(\alpha\rho+c)\lambda\sqrt{s}+4\lambda\alpha(1-\rho)\left\|\boldsymbol{\beta}^{*}\right\|_{2}}{2\lambda(1-\rho)+\phi_{\mathrm{TENet}}},$
$U_{\mathrm{ENet}}$ $=\frac{2(\rho+c)\lambda\sqrt{s}+4\lambda(1-\rho)\left\|\boldsymbol{\beta}^{*}\right\|_{2}}{2\lambda(1-\rho)+\phi_{\mathrm{ENet}}}.$
From Lemma 7, $1/\phi_{\mathrm{TENet}}\leq 1/\phi_{\mathrm{ENet}}$. Since $\alpha\leq 1$,
$2(\alpha\rho+c)\lambda\sqrt{s}+4\lambda\alpha(1-\rho)\left\|\boldsymbol{\beta}^{*}\right\|_{2}$ $\leq 2(\rho+c)\lambda\sqrt{s}+4\lambda(1-\rho)\left\|\boldsymbol{\beta}^{*}\right\|_{2}.$
Therefore, we obtain $U_{\mathrm{TENet}}\leq U_{\mathrm{ENet}}$. ∎
Appendix C Proof of Proposition 4
We prove Proposition 4 as follows.
Proof of Proposition 4.
Suppose $\tilde{\boldsymbol{\beta}}=\boldsymbol{\beta}^{*}$. Then, we have
$U_{\mathrm{TENet}}$ $=\frac{2(\alpha\rho+c)\lambda\sqrt{s}+4\lambda\alpha(1-\rho)\left\|\boldsymbol{\beta}^{*}\right\|_{2}}{2\lambda(1-\rho)+\phi_{\mathrm{TENet}}},$
$U_{\mathrm{TLasso}}$ $=\frac{2(\alpha+c)\lambda\sqrt{s}}{\phi_{\mathrm{ENet}}}.$
Regarding the numerator, from $\sqrt{s}/2\geq\|\boldsymbol{\beta}^{*}_{S}\|_{2}$,
$2(\alpha\rho+c)\lambda\sqrt{s}+4\lambda\alpha(1-\rho)\left\|\boldsymbol{\beta}^{*}\right\|_{2}$ $=2(\alpha\rho+c)\lambda\sqrt{s}+4\lambda\alpha(1-\rho)\left\|\boldsymbol{\beta}^{*}_{S}\right\|_{2}$
$\leq 2(\alpha\rho+c)\lambda\sqrt{s}+2\lambda\alpha(1-\rho)\sqrt{s}$
$=2(\alpha+c)\lambda\sqrt{s}.$
Due to $\phi_{\mathrm{TENet}}+2\lambda(1-\rho)\geq\phi_{\mathrm{TLasso}}$, we obtain $U_{\mathrm{TLasso}}\geq U_{\mathrm{TENet}}$. ∎
Appendix D Proof of Theorem 5
We establish the following proof of Theorem 5.
Proof of Theorem 5.
By differentiating $\mathcal{L}(\boldsymbol{\beta};\tilde{\boldsymbol{\beta}})$ with respect to $\beta_{j}$,
$\frac{\partial}{\partial\beta_{j}}\mathcal{L}(\boldsymbol{\beta};\tilde{\boldsymbol{\beta}})$ $:=-\frac{1}{n}\sum_{i=1}^{n}X_{i,j}\left(y_{i}-\boldsymbol{\beta}^{\top}\boldsymbol{X}_{i}\right)+\lambda\alpha\rho\operatorname{sgn}(\beta_{j})+2\lambda\alpha(1-\rho){\beta}_{j}$
$\quad+\lambda(1-\alpha)\rho\operatorname{sgn}({\beta}_{j}-\tilde{\beta}_{j})+2\lambda(1-\alpha)(1-\rho)({\beta}_{j}-\tilde{\beta}_{j}).$
From the definition of the estimator, $\frac{\partial}{\partial\beta_{j}}\mathcal{L}(\hat{\boldsymbol{\beta}};\tilde{\boldsymbol{\beta}})=0$ because of strict convexity of $\mathcal{L}(\boldsymbol{\beta};\tilde{\boldsymbol{\beta}})$ for $\rho\neq 1$. Therefore, if $\hat{{\beta}}_{j}\hat{{\beta}}_{k}>0$ and $(\hat{{\beta}}_{j}-\tilde{{\beta}}_{j})(\hat{{\beta}}_{k}-\tilde{{\beta}}_{k})>0$, then we have
$\frac{\partial}{\partial\beta_{j}}\mathcal{L}(\hat{\boldsymbol{\beta}};\tilde{\boldsymbol{\beta}})-\frac{\partial}{\partial\beta_{k}}\mathcal{L}(\hat{\boldsymbol{\beta}};\tilde{\boldsymbol{\beta}})$
$=-\frac{1}{n}\sum_{i=1}^{n}\left(X_{i,j}-X_{i,k}\right)\left(y_{i}-\hat{\boldsymbol{\beta}}^{\top}\boldsymbol{X}_{i}\right)+2\lambda(1-\rho)\left(\hat{\beta}_{j}-\hat{\beta}_{k}\right)$
$\quad-2\lambda(1-\alpha)(1-\rho)(\tilde{\beta}_{k}-\tilde{\beta}_{j})=0,$
thus we obtain
$\left|\hat{\beta}_{j}-\hat{\beta}_{k}\right|$ $=\frac{1}{2n\lambda(1-\rho)}\left|\left(\boldsymbol{X}_{\cdot,j}-\boldsymbol{X}_{\cdot,k}\right)^{\top}\left(\boldsymbol{y}-\hat{\boldsymbol{\beta}}^{\top}\mathbf{X}\right)+(1-\alpha)(\tilde{\beta}_{k}-\tilde{\beta}_{j})\right|,$
$\leq\frac{1}{2n\lambda(1-\rho)}\left\|\boldsymbol{X}_{\cdot,j}-\boldsymbol{X}_{\cdot,k}\right\|_{2}\left\|\boldsymbol{y}-\hat{\boldsymbol{\beta}}^{\top}\mathbf{X}\right\|_{2}+(1-\alpha)\left|\tilde{\beta}_{k}-\tilde{\beta}_{j}\right|,$
$=\frac{1}{2n\lambda(1-\rho)}\left\|\boldsymbol{y}-\hat{\boldsymbol{\beta}}^{\top}\mathbf{X}\right\|_{2}\sqrt{2n(1-r_{jk})}+(1-\alpha)\left|\tilde{\beta}_{k}-\tilde{\beta}_{j}\right|,$
where $\boldsymbol{X}_{\cdot,j}$ denotes the $j$th column of $\mathbf{X}$. Since $\frac{\partial}{\partial\beta_{j}}\mathcal{L}(\hat{\boldsymbol{\beta}};\tilde{\boldsymbol{\beta}})\leq\frac{\partial}{\partial\beta_{j}}\mathcal{L}({\boldsymbol{0}};\tilde{\boldsymbol{\beta}})$, we have
$\frac{1}{2n}\left\|\boldsymbol{y}-\hat{\boldsymbol{\beta}}^{\top}\mathbf{X}\right\|_{2}^{2}$
$\leq\frac{1}{2n}\left\|\boldsymbol{y}-\hat{\boldsymbol{\beta}^{\top}}\mathbf{X}\right\|_{2}^{2}+\lambda\left[\alpha\left\{\rho\|\hat{\boldsymbol{\beta}}\|_{1}+(1-\rho)\|\hat{\boldsymbol{\beta}}\|_{2}^{2}\right\}+(1-\alpha)\left\{\rho\|\hat{\boldsymbol{\beta}}-\tilde{\boldsymbol{\beta}}\|_{1}+(1-\rho)\|\hat{\boldsymbol{\beta}}-\tilde{\boldsymbol{\beta}}\|_{2}^{2}\right\}\right]$
$\leq\frac{1}{2n}\left\|\boldsymbol{y}\right\|_{2}^{2}+\lambda(1-\alpha)\rho\|\tilde{\boldsymbol{\beta}}\|_{1}+\lambda(1-\alpha)(1-\rho)\|\tilde{\boldsymbol{\beta}}\|_{2}^{2}.$
From the results above, we obtain
$\left|\hat{\beta}_{j}-\hat{\beta}_{k}\right|$ $\leq\sqrt{\frac{\left\|\boldsymbol{y}\right\|_{2}^{2}+2n\lambda(1-\alpha)\rho\|\tilde{\boldsymbol{\beta}}\|_{1}+2n\lambda(1-\alpha)(1-\rho)\|\tilde{\boldsymbol{\beta}}\|_{2}^{2}}{2n\lambda^{2}(1-\rho)^{2}}}\sqrt{1-r_{jk}}+(1-\alpha)\left|\tilde{\beta}_{k}-\tilde{\beta}_{j}\right|.$